OpenStudy (anonymous):

-1 5 -5 -- -- -- 6 3 6 List from least to greatest

OpenStudy (anonymous):

\[\frac{ -1 }{ 6},\frac{ 5 }{ 3 },\frac{ -5 }{ 6 } least \to greatest \]

OpenStudy (lgbasallote):

hint: \[\frac 53 \implies \frac {10}6\] does that help?

OpenStudy (anonymous):

Yup ty :P

OpenStudy (lgbasallote):

welcome